Find the area of the region under the graph of f on [a, b].
f(x) = x^2 − 2x + 3; [−1, 2]

Answer by Alan3354(69443)

Find the area of the region under the graph of f on [a, b].
f(x) = x^2 - 2x + 3; [-1, 2]
------------
INT = x^3/3 - x^2 + 3x + C
----
Ignore the constant C
INT(2) = 8/3 - 4 + 6 = 14/3
INT(-1) = -1/3 - 1 - 3 = -13/3
---
INT(2) - INT(-1) = 9 sq units